a car and bus leave from the same spot and drive in opposite directions. the bus is travelling 6 miles per hour slower than the car. in 4.50 hours, the car and the bus are 217 miles apart. find the average speed of the car

Answer by mananth(16946)

a car and bus leave from the same spot and drive in opposite directions. the bus is travelling 6 miles per hour slower than the car. in 4.50 hours, the car and the bus are 217 miles apart. find the average speed of the car
Let speed of car be x m/hr
Bus speed will be (x-6) m/hr
they are moving in opposite directions
d= s*t
4.5x+4.5(x-6)=217
9x-27=217
9x= 244
x=27.11
The average speed of the car is approximately 27.11 miles per hour